# A Training-Free Length Extrapolation Approach for LLMs: Greedy Attention Logit Interpolation (GALI)

### v1

@@ 4.5 Ablation Studies
In this section, we investigate the impact of local window size and chunk size on GALI. We conducted our experiments using NarrativeQA, the longest dataset in LongBench. The results are shown in Figure 5. First, we observe that the differences across the three local window sizes are minimal, indicating that attention logit interpolation effectively approximates the true attention score distribution.
Secondly, as the chunk size increases, we hypothesize that the observed effects result from the interplay of two factors. Initially, a smaller chunk size aligns better with GALI’s design, which prioritizes leveraging pretrained positional intervals as much as possible while minimizing the number of interpolations for each token. Consequently, when the chunk size increases, the number of pretrained positional intervals utilized by each token decreases, while the number of interpolated positional intervals increases, leading to performance degradation.
However, as the chunk size grows, more tokens have their positional intervals compressed into a smaller range. As analyzed earlier, performing denser interpolations within a smaller positional interval range, such as [0, 4096), yields better results than performing sparser interpolations over a larger positional interval range, such as [0, 8192). Therefore, the performance of GALI begins to improve as the chunk size further increases.
@@ 5 Conclusion
In this paper, we introduced Greedy Attention Logit Interpolation (GALI), a novel training-free length extrapolation method, and evaluated it against other approaches across widely used real-world long-context tasks, long-context language modeling, and attention distribution analysis, demonstrating its superiority. The experiment results further reveal that LLMs interpret positional intervals differently within their training context window, suggesting that length extrapolation within a smaller positional range may outperform the vanilla LLM—even for short-context tasks.
Originally designed for RoPE-based LLMs, GALI avoids direct position embedding computations, making it compatible with any LLM exhibiting long-term decay, such as ALiBi. Future work will explore its applicability to other position embedding mechanisms. However, GALI currently requires two passes of attention logit computation, introducing computational overhead and making it incompatible with flash attention. To address this, we aim to integrate interpolation estimation into flash attention and develop more efficient local interpolation techniques.
6 Impact Statement
GALI is a training-free length extrapolation method that enables LLMs to handle long-context tasks without requiring fine-tuning. It is compatible with various position embedding schemes, making it a flexible plug-and-play solution. Furthermore, we show that, in practical scenarios, interpolating within a smaller context window than the training context window can enhance performance across both long-context and short-context downstream tasks.
Appendix A Pseudo code of GALI
@@
In this section, we provide the pseudo-code for the key steps required to implement GALI. Algorithm 1 generates the chunk sizes needed to partition the input during the prefill phase. While this function can be modified to support dynamic chunk sizes, we use fixed chunk sizes in our experiments to better control memory usage. Algorithm 2 interpolates new position IDs based on the minimum number of new IDs required for each chunk. Algorithm LABEL:Interpolated_attention_logits demonstrates how we perform attention logit interpolation. Note that we use $r=\lceil{m}\rceil-n$ to represent the interval between $q_{m}$ and $k_{n}$. This is because, when computing attention logits using RoPE, we cannot directly manipulate the relative positional interval matrix; instead, we modify the relative positional interval matrix by separately operating on $query\_states$ and $key\_states$. By using $r=\lceil{m}\rceil-n$, we ensure that $\lfloor{r}\rfloor=\lceil{m}\rceil-\lceil{n}\rceil$ and $\lceil{r}\rceil=\lceil{m}\rceil-\lfloor{n}\rfloor$, enabling modifications to the relative positional interval matrix while preserving the relative order between $query\_states$ and $key\_states$. It is important to note that some operations, such as reshaping, which do not affect the core concept, are omitted from the pseudo-code in these three algorithms.
@@
0:  Interpolated attention logits # Compute the rotary embedding
@@
7:  $key\_states\_floor\leftarrow(key\_states\cdot cos\_floor)+(\text{rotate\_half}(key\_states)\cdot sin\_floor)$ Compute attention logits with $\lceil{R}\rceil and\lfloor{R}\rfloor$
@@
11:  $attn\_logits\leftarrow attn\_floor-(attn\_floor-attn\_ceil)\cdot rel\_coef$ Add normal distribution noise
@@
15:  $noise\leftarrow\text{torch.randn\_like}(attn\_logits)$
@@
18:  $attn\_logits\leftarrow attn\_logits+noise$
19:  return $attn\_logits$
Appendix B Data stastics
In this section, we provide detailed information about each dataset used in LongBench and L-Eval. Table LABEL:tab:benchmark_task_type presents the word length, task type, and number of samples for each dataset. Figure 6(a) and 6(b) show the length distributions of each dataset using the Llama2 and Llama3 tokenizers, respectively.
Benchmark Dataset Task Type Avg Len #Sample
LongBench NarrativeQA Single-doc QA 18409 200
Qasper Single-doc QA 3619 200
MultiField-en Single-doc QA 4559 150
HotpotQA Multi-doc QA 9151 200
2WikiMQA Multi-doc QA 4887 200
Musique Multi-doc QA 11214 200
GovReport Summarization 8734 200
QMSum Summarization 10614 200
MultiNews Summarization 2113 200
TREC Few shot 5177 200
TriviaQA Few shot 8209 200
SAMSum Few shot 6258 200
PassageCount Synthetic 11141 200
PassageRe Synthetic 9289 200
LCC Code 1235 500
RepoBench-P Code 4206 500
L-Eval Coursera Multiple choice 9075 172
L-Eval GSM (16-shot) Solving math problems 5557 100
QuALITY Multiple choice 7169 202
TOEFL Multiple choice 3907 269
SFCition True or False Questions 16381 64
CodeU Deducing program outputs 31575 90
(a) Llama2 Tokenizer
(b) Llama3 Tokenizer
Appendix C Implementation details
In this section, we provide detailed implementation information for each method. For Dyn-NTK and YaRN, we utilize the implementations available in Huggingface [^3] https://huggingface.co by adding rope_scaling = {”rope_type”:”dynamic”} and rope_scaling = {”rope_type”:”yarn”}, respectively, to the LLM’s config.json file. For NTK, we implement it by adding rope_scaling = {”rope_type”:”dynamic”} and static_ntk=True, and modifying the _dynamic_frequency_update function of the LlamaRotaryEmbedding class as follows:
@@ Appendix C Implementation details
self.register_buffer("inv_freq", inv_freq, persistent=False) # TODO joao: may break with compilation
@@
Length Extrapolation Method Hyperparameters
2k to 8k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 4}
@@
4k to 8k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 2}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 2}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 2}
@@
4k to 16k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 4}
@@
4k to 32k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 8}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 8}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 8}
@@
8k to 16k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 2}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 2}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 2}
@@
8k to 32k NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, ”factor”: 4}
YaRN rope_scaling={”rope type”:”YaRN”, ”factor”: 4}
@@
Appendix D Extra experiment results
D.1 Real-world long-context task results
Regarding the best-performing method on Llama2-4k, SelfExtend has been reported to be highly sensitive to hyperparameters (Jin et al., 2024). Specifically, larger group sizes and smaller local windows sometimes yield better results, which supports our conclusion in Section 4.2. These configurations emphasize the use of smaller positional intervals, reducing reliance on larger ones and preventing content from being placed in less well-understood positional intervals. This limitation affects GALI’s effectiveness on Llama2, as GALI assumes the model fully understands its entire training context window, thereby always maximizing the use of pretrained positional intervals.
NarrativeQA
Qasper
MultiField-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PassageCount
PassageRe
Lcc
RepoBench-P
Table 6: Performance comparison with different backbone LLMs and training-free length extrapolation methods. The best result in each experiment has been bolded. * indicates the results reported by LongBench(Bai et al., 2024), * indicates the results reported by LongBench(Jin et al., 2024). The number following each method represents the target context window. For example, 16k means 16 × 1024. The ”Original” means testing with the backbone model, i.e., the model shown in the left column.
 Methods Coursera GSM QuALITY TOFEL SFiction CodeU Average
Original 29.80 29.00 37.62 58.36 60.16 1.11 36.01
SelfExtend-16k* 35.76 25.00 41.09 55.39 57.81 1.11 36.02
SelfExtend-16k 32.99 29.00 40.59 57.62 57.81 2.22 36.71
ChunkLlama* 32.12 31.00 35.14 57.62 61.72 2.22 36.64
ChunkLlama-16k 28.92 31.00 43.07 58.36 60.94 2.22 37.42
NTK-16k* 32.71 19.00 33.16 52.78 64.84 0.00 33.75
NTK-16k 26.89 16.00 33.66 60.97 41.41 0.00 29.82
Dyn-NTK* 13.95 13.00 30.69 52.27 57.02 1.11 28.01
Dyn-NTK-16k 15.41 13.00 33.17 54.65 54.69 1.11 28.67
YaRN-16k 36.49 18.00 42.08 57.62 42.97 7.78 34.15
(Ours)GALI-16k 35.32 29.00 39.11 54.65 51.43 4.44 35.66
Llama3-8b-ins-8k Original 53.05 58.00 61.88 82.16 60.16 4.44 53.93
SelfExtend-16k 55.38 63.00 62.87 82.16 64.06 5.56 55.76
ChunkLlama* 56.24 54.00 63.86 83.27 70.31 5.56 55.54
ChunkLlama-16k 53.34 54.00 60.89 81.78 61.72 5.56 53.53
NTK-16k 52.03 77.00 65.35 81.04 42.97 0.00 56.58
Dyn-NTK-16k 52.03 55.00 61.88 82.16 52.34 2.22 52.89
YaRN-16k 55.96 75.00 63.37 79.93 62.50 5.56 57.83
(Ours)GALI-16k 54.65 59.09 61.88 83.33 65.63 6.67 55.42
ChunkLlama-32k 54.36 55.00 60.89 81.78 64.06 5.56 53.61
NTK-32k 58.28 83.00 63.86 81.04 59.38 1.11 57.78
Dyn-NTK-32k 54.36 55.00 61.88 82.16 64.06 6.67 54.02
YaRN-32k 55.23 76.00 62.38 79.18 67.19 7.78 57.96
(Ours)GALI-32k 54.17 59.09 62.38 82.68 66.41 7.78 55.29
Table 7: Performance comparison with different backbone LLMs and training-free length extrapolation methods. The best result in each experiment has been bolded. * indicates the results reported by ChunkLlama(An et al., 2024b), * indicates the results reported by LongBench(Jin et al., 2024), and * indicates the results reported by L-Eval(An et al., 2024a).
D.2 Long language modeling task results
Methods 1k 2k 3k 4k 5k 6k 7k 8k
Llama3-7b-chat-4k SelfExtend 8.81 8.99 9.16 9.24 9.25 9.16 9.2 9.3
ChunkLlama 9.07 9.26 9.41 9.45 9.43 9.31 9.31 9.39
NTK 8.95 9.04 9.16 9.18 9.16 9.06 9.26 13.67
Dyn-NTK 8.81 8.99 9.15 10.79 44.32 87.35 160.07 224.07
YaRN 11.65 8.97 9.07 9.16 9.17 9.15 9.03 9.04
(Ours)GALI 8.81 8.99 9.15 9.24 9.59 9.66 9.63 9.66
D.3 Attention distribution analysis results
In this section, we present the detailed results of the attention distribution analysis. First, we compare the differences between the attention score matrix of length extrapolation methods and the standard attention score matrix, as shown in Figure LABEL:attn_score_dist. For this analysis, we averaged the attention score matrices for each layer and each head before comparison. Whether comparing Llama3-2k or Llama3-4k, GALI consistently achieved the highest similarity to the standard attention score matrix. Additionally, we observed that all methods exhibited higher values in the lower-left corner of the matrix compared to the standard attention score matrix. We attribute this to the fact that these methods do not perform true extrapolation, whereas the standard Llama3-8k model, utilizing a larger positional interval range [0, 8192), results in a lower mean value of the attention scores.
(a) Llama3-8b-ins-2k backbone
(b) Llama3-8b-ins-4k backbone
Figure 7: This is a comparison of the attention score matrices obtained using Llama3-2k and Llama3-4k for length extrapolation with those of Llama3-8k. Note that we averaged the attention scores across all layers and heads, applied average pooling to scale the matrix to 0.05%, and set the maximum value of the heatmap to 0.0005 for better visualization. “Original” represents the attention score matrix of Llama3-8k, and the number next to each method’s name indicates the sum of the absolute differences between the method’s attention scores matrices and the “Original” matrix.
(a) Attention scores of row 1000
(b) Attention scores of row 2000
(c) Attention scores of row 3000
(d) Attention scores of row 4000
(e) Attention scores of row 5000
(f) Attention scores of row 6000
(g) Attention scores of row 7000
(h) Attention scores of row 8000
(a) Attention scores of row 1000
(b) Attention scores of row 2000
(c) Attention scores of row 3000
(d) Attention scores of row 4000
(e) Attention scores of row 5000
(f) Attention scores of row 6000
(g) Attention scores of row 7000
(h) Attention scores of row 8000

### v2

@@ 4.5 Ablation Studies
In this section, we investigate the impact of the size of the local window and the chunk on GALI. We conducted our experiments using NarrativeQA, the longest dataset in LongBench. The results are in Figure 4. First, we observe that the differences across the three local window sizes are marginal, indicating that attention logit interpolation effectively approximates the true attention score distribution. Secondly, as the chunk size increases, we hypothesize that the observed effects result from the interplay of two factors. Initially, a smaller size of the chunk aligns better with the design of GALI, which prioritizes leveraging pretrained positional intervals as much as possible while minimizing the number of interpolations for each token. When the chunk size increases, the number of pretrained positional intervals utilized by each token decreases, while the number of interpolated positional intervals increases, leading to performance degradation. However, as the chunk size grows, more tokens have their positional intervals compressed into a smaller range. As analyzed earlier, performing denser interpolations within a smaller positional interval range, such as [0, 4096), yields better results than performing sparser interpolations over a larger positional interval range, such as [0, 8192). GALI’s performance begins to improve.
We also performed an ablation study on the use of Gaussian noise to assess its impact. In Table 4, the model exhibited a slight performance degradation when Gaussian noise was removed, suggesting that simulating oscillatory behavior via Gaussian perturbation is indeed beneficial. Nevertheless, due to the overall downward trend of attention logit over long sequences, attention logit interpolation remains effective even in the absence of noise.
Table 4: Noise Analysis on LongBench and L-Eval.
Llama3-8b-ins-8k Model Noise Average
GALI-16k No 44.45
GALI-16k Yes 45.17
GALI-32k No 44.53
GALI-32k Yes 45.38
(a) LongBench results
Llama3-8b-ins-8k Model Noise Average
GALI-16k No 42.01
GALI-16k Yes 42.32
GALI-32k No 40.16
GALI-32k Yes 42.79
(b) L-Eval results
@@ 5 Conclusion
The paper introduces Greedy Attention Logit Interpolation (GALI), a training-free method for length extrapolation in LLMs. Our evaluations show GALI achieves stable and superior performance across both short- and long-context tasks without requiring any input-length-specific tuning. We found that extrapolation within narrower positional ranges can yield better results, even on short context tasks. GALI avoids computation over position embeddings, making it compatible with other architectures exhibiting long-term decay, such as ALiBi.
Limitations
GALI’s current limitation is the need for two passes of attention logit computation, making it incompatible with flash attention. Future work will focus on integrating GALI into flash attention and improving the efficiency of local linear interpolation.
Appendix A Long term decay of RoPE
We show another example of the long term decay caused by RoPE in this section.
Figure 5: Visualization of long-term decay in attention logit. The sentence “The quick brown fox jumps over the lazy dog.” is fed into a one-layer Llama3-8b-ins model. Figure 5 shows the attention logit from the last token to the second-to-last token. Another example can be found in Appendix Figure 5 presents the logit from the last token to the first token. As the positional ID interval increases from 1 to 819, a clear decay phenomenon in the logit is observed.
Appendix B Data stastics
In this section, we provide detailed information about each dataset used in LongBench and L-Eval. Table 5 presents the word length, task type, and number of samples for each dataset. Figure 6(a) and 6(b) show the length distributions of each dataset using the Llama2 and Llama3 tokenizers, respectively.
Benchmark Dataset Task Type Avg Len #Sample
LongBench NarrativeQA Single-doc QA 18409 200
Qasper Single-doc QA 3619 200
MultiField-en Single-doc QA 4559 150
HotpotQA Multi-doc QA 9151 200
2WikiMQA Multi-doc QA 4887 200
Musique Multi-doc QA 11214 200
GovReport Summarization 8734 200
QMSum Summarization 10614 200
MultiNews Summarization 2113 200
TREC Few shot 5177 200
TriviaQA Few shot 8209 200
SAMSum Few shot 6258 200
PassageCount Synthetic 11141 200
PassageRe Synthetic 9289 200
LCC Code 1235 500
RepoBench-P Code 4206 500
L-Eval Coursera Multiple choice 9075 172
GSM (16-shot) Solving math problems 5557 100
QuALITY Multiple choice 7169 202
TOEFL Multiple choice 3907 269
SFCition True or False Questions 16381 64
CodeU Deducing program outputs 31575 90
(a) Llama2 Tokenizer
(b) Llama3 Tokenizer
Appendix C Extra experiment results
C.1 Real-world long-context task results
Regarding the best-performing method on Llama2-4k, SelfExtend has been reported to be highly sensitive to hyperparameters Jin et al. (2024). Specifically, larger group sizes and smaller local windows sometimes yield better results, which supports our conclusion in Section 4.2. These configurations emphasize the use of smaller positional intervals, reducing reliance on larger ones and preventing content from being placed in less well-understood positional intervals. This limitation affects GALI’s effectiveness on Llama2, as GALI assumes the model fully understands its entire training context window, thereby always maximizing the use of pretrained positional intervals.
Table 6: Performance comparison with different backbone LLMs and training-free length extrapolation methods. The best result in each experiment has been bolded. * indicates the results reported by LongBenchBai et al. (2024), * indicates the results reported by LongBenchJin et al. (2024). The number following each method represents the target context window. For example, 16k means 16 × 1024. The "Original" means testing with the backbone model, i.e., the model shown in the left column.
NarrativeQA
Qasper
MultiField-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PassageCount
PassageRe
Lcc
RepoBench-P
Table 7: Performance comparison with different backbone LLMs and training-free length extrapolation methods. The best result in each experiment has been bolded. * indicates the results reported by ChunkLlamaAn et al. (2024b), * indicates the results reported by LongBenchJin et al. (2024), and * indicates the results reported by L-EvalAn et al. (2024a).
 Methods Coursera GSM QuALITY TOFEL SFiction CodeU Average
Original 29.80 29.00 37.62 58.36 60.16 1.11 36.01
SelfExtend-16k* 35.76 25.00 41.09 55.39 57.81 1.11 36.02
SelfExtend-16k 32.99 29.00 40.59 57.62 57.81 2.22 36.71
ChunkLlama* 32.12 31.00 35.14 57.62 61.72 2.22 36.64
ChunkLlama-16k 28.92 31.00 43.07 58.36 60.94 2.22 37.42
NTK-16k* 32.71 19.00 33.16 52.78 64.84 0.00 33.75
NTK-16k 26.89 16.00 33.66 60.97 41.41 0.00 29.82
Dyn-NTK* 13.95 13.00 30.69 52.27 57.02 1.11 28.01
Dyn-NTK-16k 15.41 13.00 33.17 54.65 54.69 1.11 28.67
YaRN-16k 36.49 18.00 42.08 57.62 42.97 7.78 34.15
(Ours)GALI-16k 35.32 29.00 39.11 54.65 51.43 4.44 35.66
Llama3-8b-ins-8k Original 53.05 58.00 61.88 82.16 60.16 4.44 53.93
SelfExtend-16k 55.38 63.00 62.87 82.16 64.06 5.56 55.76
ChunkLlama* 56.24 54.00 63.86 83.27 70.31 5.56 55.54
ChunkLlama-16k 53.34 54.00 60.89 81.78 61.72 5.56 53.53
NTK-16k 52.03 77.00 65.35 81.04 42.97 0.00 56.58
Dyn-NTK-16k 52.03 55.00 61.88 82.16 52.34 2.22 52.89
YaRN-16k 55.96 75.00 63.37 79.93 62.50 5.56 57.83
(Ours)GALI-16k 54.65 59.09 61.88 83.33 65.63 6.67 55.42
ChunkLlama-32k 54.36 55.00 60.89 81.78 64.06 5.56 53.61
NTK-32k 58.28 83.00 63.86 81.04 59.38 1.11 57.78
Dyn-NTK-32k 54.36 55.00 61.88 82.16 64.06 6.67 54.02
YaRN-32k 55.23 76.00 62.38 79.18 67.19 7.78 57.96
(Ours)GALI-32k 54.17 59.09 62.38 82.68 66.41 7.78 55.29
C.2 Long language modeling task results
Methods 1k 2k 3k 4k 5k 6k 7k 8k
Llama3-7b-chat-4k SelfExtend 8.81 8.99 9.16 9.24 9.25 9.16 9.2 9.3
ChunkLlama 9.07 9.26 9.41 9.45 9.43 9.31 9.31 9.39
NTK 8.95 9.04 9.16 9.18 9.16 9.06 9.26 13.67
Dyn-NTK 8.81 8.99 9.15 10.79 44.32 87.35 160.07 224.07
YaRN 11.65 8.97 9.07 9.16 9.17 9.15 9.03 9.04
(Ours)GALI 8.81 8.99 9.15 9.24 9.59 9.66 9.63 9.66
C.3 Attention distribution analysis results
In this section, we present the detailed results of the attention distribution analysis. First, we compare the differences between the attention score matrix of length extrapolation methods and the standard attention score matrix, as shown in Figure 7. For this analysis, we averaged the attention score matrices for each layer and each head before comparison. Whether comparing Llama3-2k or Llama3-4k, GALI consistently achieved the highest similarity to the standard attention score matrix. Additionally, we observed that all methods exhibited higher values in the lower-left corner of the matrix compared to the standard attention score matrix. We attribute this to the fact that these methods do not perform true extrapolation, whereas the standard Llama3-8k model, utilizing a larger positional interval range [0, 8192), results in a lower mean value of the attention scores.
(a) Llama3-8b-ins-2k backbone
(b) Llama3-8b-ins-4k backbone
Figure 7: This is a comparison of the attention score matrices obtained using Llama3-2k and Llama3-4k for length extrapolation with those of Llama3-8k. Note that we averaged the attention scores across all layers and heads, applied average pooling to scale the matrix to 0.05%, and set the maximum value of the heatmap to 0.0005 for better visualization. “Original” represents the attention score matrix of Llama3-8k, and the number next to each method’s name indicates the sum of the absolute differences between the method’s attention score matrix and the “Original” matrix.
(a) Attention scores of row 1000
(b) Attention scores of row 2000
(c) Attention scores of row 3000
(d) Attention scores of row 4000
(e) Attention scores of row 5000
(f) Attention scores of row 6000
(g) Attention scores of row 7000
(h) Attention scores of row 8000
(a) Attention scores of row 1000
(b) Attention scores of row 2000
(c) Attention scores of row 3000
(d) Attention scores of row 4000
(e) Attention scores of row 5000
(f) Attention scores of row 6000
(g) Attention scores of row 7000
(h) Attention scores of row 8000
Appendix D Pseudo code of GALI
@@
In this section, we provide the pseudo-code for the key steps required to implement GALI. Algorithm 1 generates the chunk sizes needed to partition the input during the prefill phase. While this function can be modified to support dynamic chunk sizes, we use fixed chunk sizes in our experiments to better control memory usage. Algorithm 2 interpolates new position IDs based on the minimum number of new IDs required for each chunk. Algorithm 3 demonstrates how we perform attention logit interpolation. Note that we use $r=\lceil{m}\rceil-n$ to represent the interval between $q_{m}$ and $k_{n}$. This is because, when computing attention logit using RoPE, we cannot directly manipulate the relative positional interval matrix; instead, we modify the relative positional interval matrix by separately operating on $query\_states$ and $key\_states$. By using $r=\lceil{m}\rceil-n$, we ensure that $\lfloor{r}\rfloor=\lceil{m}\rceil-\lceil{n}\rceil$ and $\lceil{r}\rceil=\lceil{m}\rceil-\lfloor{n}\rfloor$, enabling modifications to the relative positional interval matrix while preserving the relative order between $query\_states$ and $key\_states$. It is important to note that some operations, such as reshaping, which do not affect the core concept, are omitted from the pseudo-code in these three algorithms.
@@
0:  Interpolated attention logit # Compute the rotary embedding
@@
7:  $key\_states\_floor\leftarrow(key\_states\cdot cos\_floor)+(\text{rotate\_half}(key\_states)\cdot sin\_floor)$ Compute attention logit with $\lceil{R}\rceil and\lfloor{R}\rfloor$
@@
11:  $attn\_logit\leftarrow attn\_floor-(attn\_floor-attn\_ceil)\cdot rel\_coef$ Add normal distribution noise
@@
15:  $noise\leftarrow\text{torch.randn\_like}(attn\_logit)$
@@
18:  $attn\_logit\leftarrow attn\_logit+noise$
19:  return $attn\_logit$
Appendix E Implementation details
In this section, we provide detailed implementation information for each method. For Dyn-NTK and YaRN, we utilize the implementations available in Huggingface [^3] https://huggingface.co by adding rope_scaling = {"rope_type":"dynamic"} and rope_scaling = {"rope_type":"yarn"}, respectively, to the LLM’s config.json file. For NTK, we implement it by adding rope_scaling = {"rope_type":"dynamic"} and static_ntk=True, and modifying the _dynamic_frequency_update function of the LlamaRotaryEmbedding class as shown in Table 9.
Table 9: The implementation of NTK used in our experiments.
@@ Appendix E Implementation details
self.register_buffer("inv_freq", inv_freq, persistent=False)
@@
Exp. Method Hyperparameters
2k to 8k NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 4}
@@
4k to 8k NTK rope_scaling={”rope type”:”dynamic”, "factor": 2}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 2}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 2}
@@
4k to 16k NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 4}
@@
4k to 32k NTK rope_scaling={”rope type”:”dynamic”, "factor": 8}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 8}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 8}
@@
8k to 16k NTK rope_scaling={”rope type”:”dynamic”, "factor": 2}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 2}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 2}
@@
8k to 32k NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}, static_ntk=True
Dyn-NTK rope_scaling={”rope type”:”dynamic”, "factor": 4}
YaRN rope_scaling={”rope type”:"YaRN", "factor": 4}